OpenStudy (anonymous):

Medals being given! the rectangular coordinates of a point given. Find the polar coordinates of each point! 1. (4,0) 2. (3,4) 3. (2,-2)

hartnn (hartnn):

\(\Large r =\sqrt{x^2+y^2 } \\ \Large \theta = tan^{-1} (\dfrac{y}{x})\)

OpenStudy (anonymous):

okay, got them! Im disappointed my notes didnt contain the formula to do this. Thank you again!! :)

hartnn (hartnn):

welcome ^_^ internet is the source of infinite information!

OpenStudy (anonymous):

yeah, but most of the math sites im finding are practically molded with super fancy math lingo it takes a long time to decipher. So if you can suggest any you know specifically, that would be really great!

hartnn (hartnn):

i suggest \(\Large OpenStudy\) :P

hartnn (hartnn):

'mathisfun' is also too good :)

OpenStudy (anonymous):

nice! XD and thank you! im checking it out now